# Anatomy of Elite and Mass Polarization in Social Networks

### v1

@@ 2.2 Defining structural polarization and issue alignment for hierarchies
Prior work has shown that the most common quantitative methods for measuring structural polarization in social networks perform similarly after they are denoised [SCK22]. According to the same work, the best performing method in detecting polarized networks was Adaptive EI-index (AEI), which is based on the more popular EI-index (e.g. [EDVBP22, Yan16, DVB18, Bor23]) developed by Krackhardt and Stern [KS88]. This method produces a score based on the distribution of connections among nodes within and between groups, where a greater density of links within groups relative to those between groups indicates a more polarized system.
@@
In the AEI-index, all links are treated homogeneously within each group, which prevents it from distinguishing the activity driven by the elite versus the mass. However, one should not view these links equal as the elite are often seen as highly influential individuals on social platforms due to their large following, prominent status in society and access to resources that can amplify their messages and shape the public opinion. An elite individual can also connect to other elite members enabling them to reach even a wider audience and have a greater impact on conversations and trends within their communities, making them “more organized”. The amplification and social validation come mostly from the followers, the surrounding mass, who both take cues from the elite and can be mobilized by them to take specific actions. Therefore, we want to distinguish the link densities according to the hierarchies in the network. We do this by decomposing the elements in $P_{AEI}$ further
@@
where $I_{c_{X}}$ represents the interaction between group $X$’s core nodes, while $I_{cp_{X}}$ denotes the interaction between its core and periphery, $I_{p_{X}}$ denotes the interaction among its periphery nodes, and $n_{X}$ denotes the number of nodes in group $X$. We can now subsitute the corresponding terms of Eq. (2) into Eq. (1), which gives us
@@
$P_{AEI}=(\underbrace{\frac{i_{c_{A}}}{\alpha}}_{\text{elite cohesion}}+\underbrace{\frac{i_{cp_{A}}}{\alpha}}_{\text{mass amplification}}+\underbrace{\frac{i_{p_{A}}}{\alpha}}_{\text{mass cohesion}})+(\frac{i_{c_{B}}}{\alpha}+\frac{i_{cp_{B}}}{\alpha}+\frac{i_{p_{B}}}{\alpha})-\underbrace{\frac{2\times e_{AB}}{\alpha}}_{\text{bridge}}\,,$ (3)
where $\alpha$ is the normalization factor equal to $(i_{c_{A}}+i_{cp_{A}}+i_{p_{A}})+(i_{c_{B}}+i_{cp_{B}}+i_{p_{B}})+2\times e_{AB}$.
@@
We call this decomposition polarization spectrum as it allows us to further study the direct effects of different groups and hierarchies on the polarization score. By splitting the observed polarization score into components, we can detect the differences between groups and their hierarchies on the polarization measure and determine whether their contributions have changed over time. When we say ’contribution’ or ’impact’ of a certain group on observed polarization, we specifically refer to these components. This is valuable information in the context of polarization, where we are often interested in monitoring the evolution of the opposing groups and their power dynamics.
Introducing hierarchical groups also enables us to operationalize the polarization processes. We call $i_{c_{X}}/\alpha$ the elite cohesion of group $X$. The higher the value, the more organized is the core. Similarly for $i_{p_{X}}/\alpha$, which we call mass cohesion of group $X$. Structurally, this value quantifies the centralization of leadership in a given group. Lower values indicate more centralized leadership, as most links would be directed towards amplifying the core rather than connecting with peers in the periphery. Finally, we name $i_{cp_{X}}/\alpha$ as mass amplification, capturing the intuition that surrounding periphery act as an amplifying body for its group’s core voices.
@@
We expect the hierarchy of users playing a significant role in terms of polarization. To better understand the impact on score, we also measure the marginal polarization for both core and periphery nodes. We define marginal polarization as the change in structural polarization when a node of a specific type is added to the network, which enables us to assess how introducing an average elite user versus an average non-elite user affects the network’s overall polarization.
@@
A core node in group $A$ has $k_{c_{A}}$ links to other core nodes in the same group, it receives $k_{cp_{A}}$ links from the periphery and forms $k_{out}$ external links to group $B$. For a periphery node belonging to group $A$, it can have $k_{p_{A}}$ links to other periphery nodes. It amplifies $k_{cp_{A}}$ core nodes and connects to $k_{out}$ links outside its own group. By substituting the average of observed values into these “decomposed” degrees, we can approximate the marginal polarization for a mean core node as
@@
$\Delta_{core_{A}}P_{AEI}\approx 2\times\frac{\frac{\langle k_{c_{A}}\rangle+\langle k_{cp_{A}}\rangle}{(n_{A})^{2}}-\frac{\langle k_{out_{B}}\rangle}{n_{A}n_{B}}}{\alpha}\,,$ (4)
@@
where $n_{X}$ denotes the number of nodes in group $X$. The formula is equivalent for the situation where an average periphery node is added to the polarized system $(\Delta_{periphery_{A}}P_{AEI}$), but you need to substitute the $k$ values with those reflecting the link distribution of peripheral nodes. Interestingly, the formula suggests that having an equal number of links in and out is not enough for having zero net impact on polarization when group $A$ is smaller than group $B$. See A.6 for full derivation of this marginal polarization formula.
@@
In addition to structural polarization, we aim to quantify the degree of issue alignment separately for the elite and mass. Issue alignment contributes to political polarization by creating a situation where individuals or groups align themselves strictly along specific issues. When people strongly identify with a particular issue or group and use it as a defining factor for their political stance towards other topics, it can lead to several detrimental consequences [CSG+21]. We measure alignment using the normalized mutual information (NMI) method [CSG+21] for cores and peripheries separately. In essence, it quantifies the extent to which knowing the group of an user in one topic informs us about the users’s group in another topic.
@@
Let $G_{1}=(V_{1},M_{1})$ and $G_{2}=(V_{2},M_{2})$ denote networks representing issues 1 and 2, respectively. For estimating each user’s stance on a specfic issue, we use their group membership in the given network. Since this assessment requires user’s presence in both networks, we add the stances for each node in $V_{1}\cap V_{2}$ in stance vectors, $s_{1}$ and $s_{2}$, corresponding to the respective issues, i.e. these stance vectors effectively encode each user’s group membership in the networks. Finally, we compute the issue alignment separately for the distinct hierarchies; NMI($s_{1}^{c}$, $s_{2}^{c}$) for the cores (elite alignment) and NMI($s_{1}^{p}$, $s_{2}^{p}$) for the peripheries (mass alignment). This measure is constrained between 0 and 1, where 0 indicates no alignment, and 1 indicates complete alignment between the issues in question.
@@ 3 Hierarchical polarization in Finnish Parliamentary Elections 2019 and 2023
Topic $|N|_{19}$ $|E|_{19}$ $|N|_{23}$ $|E|_{23}$ $AEI_{19}^{*}$ $AEI_{23}^{*}$
Climate politics 16 639 50 605 11 264 34 509 0.45 0.53
Immigration politics 6 739 20 923 9 519 36 889 0.50 0.61
Social policies 13 926 39 778 11 395 34 173 0.36 0.49
Economic policy 5 880 10 962 10 031 37 111 0.28 0.57
Education policy 9 843 20 314 12 814 35 619 0.29 0.47
Table 1: Retweet networks were constructed for both election years, encompassing five distinct networks each. These networks were constructed based on Twitter data obtained over a span of 12 weeks leading up to the respective election day, which were 14.4. for 2019 and 2.4. for 2023. $|N|$ represents the count of unique nodes (users), and $|E|$ denotes the count of unique edges (retweets) in network after the preprocessing. $AEI^{*}$ denotes the normalized structural polarization computed according to [SCK22], where AEI stands for Adaptive EI-index. Last column of the table shows how polarization has increased significantly for all topics studied here.
@@ 3.1 Data
We use data collected from Twitter (Public API v1 and v2) during the Finnish parliamentary elections in 2019 and 2023. To ensure consistency when comparing these two snapshots, we consider a 12-week period leading up to the election day, resulting in the following periods: from January 21, 2019, to April 14, 2019 for the first elections, and from January 9, 2023, to April 2, 2023 for the second elections. For each year, five networks are constructed from sets of keywords related to larger topics, such as immigration and climate change (see Appendix in [CSG+21] for more details). Before constructing the networks, retweets that contain keywords belonging to two or more topics are removed from the dataset, as they would produce a systematic error, i.e. inflate the extent of alignment between networks.
@@
The largest connected component is extracted for each network, as done in [GMGM18, CSG+21, SCK22], with the subsequent removal of self-loops and parallel edges, as performed in [SCK22, CSG+21]. Finally, a network can be viewed as a graph where each user contributing to the topic is assigned to one node. In this graph, an edge between two nodes represents a retweet, which can be deemed as an agreement or a shared point of view on a selected issue between the corresponding users [GMGM18, CSG+21, SCK22]. Basic information about each network can be found in Table 1, and a summary of the network partition outputs in Table 2.
@@
We tested the assumptions stated in the previous section independently: To verify the presence of political figures within cores, we assessed the probability of a politician being found in the core compared to the periphery. We found that politicians who ran for elections are at least 3-4 times more likely to be in the core than in the periphery (see A.1 for more details). To view the connections between the outer periphery and the core nodes as mass amplification, we confirmed this assumption by checking that the majority of retweets flow towards the core nodes (see A.2 for more details). Thus, we find our assumptions reasonable.
Table 1 demonstrates the drastic increase in structural polarization for all networks from 2019 to 2023. Over these four years, normalized AEI more than doubled for the economy network. The climate network showed the smallest percentage increase in polarization, approximately 18 %. For the immigration network, AEI increased from 0.50 to 0.61, reaching the highest polarization value observed among our networks. This extent of development in polarization over such a short period made us investigate its structure in more detail.
Assortative Hierarchy
Topic $G_{19}$ $G_{23}$ $A_{19}$ $B_{19}$ $A_{23}$ $B_{23}$
Climate politics
Immigration politics
Social policies
Economic policy
Education policy
Table 2: Each network studied here has an assortative structure and each identified polarized group has some type of hierarchy that could not be explained with Erdős–Rényi model. We confirm assortativity if the two-group model yields a lower description length than the single-group model. Similarly, hierarchy within each group is confirmed if the CP-model results in a lower description length compared to the null model.
@@ 3.2 Groups have unequal impact on the overall structural polarization
Figure 3: Polarization spectrum for structural contributions of different groups and their hierarchies to AEI-score. The figure illustrates the predominant influence of elite cohesion ($i_{{c}_{A}}$ $i_{{c}_{B}}$), mass amplification ($i_{{cp}_{A}}$ $i_{{cp}_{B}}$), and mass cohesion ($i_{{p}_{A}}$ $i_{{p}_{B}}$) on the overall score. The green part of spectrum represents the impact of the bridge between the opposing entities ($2\times e_{AB}$). The contributions of different hierarchical members not only vary within individual networks but also across the distinct networks. The part of the spectrum that corresponds to the internal structures is shifted to the left by an amount equal to the cross-interactions. This enables us to read the unnormalized AEI score for each network directly from the figure.
We encapsulated the contributions of different components into a single plot (see Figure 3), where separate spectrums are composed for each year separately for all topics. Here, groups labeled A (in red) refer to left-leaning groups, while groups labeled B (in blue) refer to right-leaning groups, respectively (see A.3 for labeling details). This allows us to track the changes in the dynamics of left-right groups and their embedded hierarchies in the polarized milieu online.
@@
The polarization spectrums also reveal the influence and evolution of distinct hierarchical groups. For instance, in the immigration network, the most polarized topic in both years, the elite cohesion of the right-leaning group had a relatively large impact on polarization. The same elite were also significantly more organized than those in the left-leaning group, although the elite cohesion of the latter increased in 2023. The opposite phenomenon was observed for economic policy network, where the left-leaning group’s elite cohesion was higher, while right-leaning group elite evolved into more organized structure in 2023. Note, how in all networks, at least one of the elite groups had become more cohesive during the four years; Right-leaning elite became more cohesive in four networks and left-leaning elite in three networks, indicating a stronger role of elite in certain topics. The most drastic swap in terms of contributions of a hierarchical groups was witnessed in the mass amplification patterns in climate network, where right-leaning public outrun the previously dominating left-leaning public. Although in the absolute terms, the publics’ interaction patterns tend to explain most of the observed polarization in all networks, the relative contribution (when compared to the sizes of these hierarchical groups) of the elite members is, on average, greater, which we will see later in Figures 4 & 5B.
@@
Communication between polarized groups has mostly dropped or remained at the same level across all topics. This type of communication, labeled as bridge, is the only interaction that could structurally decrease polarization. Its share of the spectrum has shrunk in all networks except the immigration network.
To summarize the changes in power dynamics between the two groups, we visualize the polarization-increasing components of AEI alongside the size distributions of groups in a more compact format (see Figure 4). For each network and year, we plot a thicker rectangle to illustrate the fraction each group (including their hierarchical groups) contributed to the observed AEI. A thinner bar shows the relative sizes of each group. The lower rectangles correspond to the data from 2019, while the upper rectangles represent the results from 2023. These rectangles, which depict the distribution of contributions, sum up to one. In a balanced scenario, the red and blue shaded areas would be symmetrical along the vertical line.
In the context of immigration, the overall impact of the right-leaning group exceeded that of the left-leaning group in both years, despite the right-leaning group being smaller in size. A drastic change was observed in the climate network, where in 2019, the left-leaning group was more dominant. However, after four years, the right-leaning group managed to organize its elite and public components more effectively, resulting in a greater impact on polarization, despite left-leaning group being larger in 2023. The plot for the economy network also illustrates how the size of a group does not necessarily correlate with its impact on the polarization score. In both years, the right-leaning groups were significantly larger, yet the left-leaning groups accounted for a larger portion of the observed polarization.
@@
Three main findings from the evolution graphs (Fig. 4) stand out: first, the increase in structural polarization does not follow uniform patterns, second, a larger group size does not always lead to a higher contribution to the observed polarization, and third, the elite play relatively higher role than the mass. The networks shown are the most polarized ones, yet the underlying contributions of the groups are very different. We see, for instance, how right-leaning group compensated its smaller size with more organized core and greater mass amplification in both years, leading to significantly higher impact on polarization when compared to the impact of left-leaning group. The greater relative contribution of the elite members is generally evident across all networks when reflecting it to the size of their groups shown in the thinner bars. We also confirm this by computing the marginal polarization separately for the elite and mass individuals in Figure 5.
Figure 4: Evolution of power structures in most polarized networks: Contributions of different structures to overall polarization is shown in the thicker bars for the most polarized networks. Mass amplification and cohesion are combined here into one color. Thinner bars illustrate the size distributions of both elite and mass groups within the networks. The upper set of bars corresponds to the results for 2023, while the lower set represents those for 2019. Each bar sums up to 1. These illustrations demonstrate visually how the increases in structural polarization does not follow uniform patterns, larger group sizes do not always lead to higher contributions to observed polarization, and the elite play a relatively higher role than the masses. The similar plots for the remaining topics can be found in Appendix A.5.
3.3 Elite is significantly more aligned than the mass
Figure 5: A: Elites are consistenly more aligned than masses across all topic pairs. Elites became more aligned in 2023, together with a smaller increase in the alignment of mass opinion on various issues. To capture the uncertainty around the observed values, we bootstrapped 500 pairs of networks for each topic pair. Each bootstrap sample represents a perturbed version of the original network, where the sizes of cores and peripheries are subject to random fluctuations. B: Elite tend to have higher marginal polarization as well compared to the mass. In both years, adding a new elite member to the economy network had the greatest impact. A weighted average of both groups’ marginal values is applied to obtain a single value representing the hierarchy’s mean effect on AEI.

### v2

@@ 3.4 Defining Structural Polarization and Issue Alignment for Groups and Hierarchies
Prior work has shown that the most common quantitative methods for measuring structural polarization in social networks perform similarly after they are denoised [56]. According to the same work, the best performing method in detecting polarized networks was Adaptive EI-index (AEI), which is based on the more popular EI-index (e.g. [17, 64, 15, 6]) developed by Krackhardt and Stern [37]. This method produces a score based on the distribution of connections among nodes within and between groups, where a greater density of links within groups relative to those between groups indicates a more polarized system.
@@
In AEI, all links are treated homogeneously within each group, which prevents it from distinguishing the activity driven by the elites versus the masses. However, one should not view these links equally as the elites are often seen as highly influential individuals on social platforms due to their large following, prominent status in society and access to resources that can amplify their messages and shape public opinion. An elite individual can also connect to other elite members enabling them to reach even a wider audience and have a greater impact on conversations and trends within their communities, making them more cohesive. The amplification and social validation come mainly from the followers, i.e., the surrounding mass, who both take cues from the elites and can be mobilized by them to take specific actions. Therefore, we want to distinguish the link densities according to the hierarchies in the network. We do this by decomposing the elements in $P_{AEI}$ further, such that
@@
where $I_{c_{X}}$ represents the interaction between group $X$’s core nodes, while $I_{cp_{X}}$ denotes the interaction between its core and periphery, $I_{p_{X}}$ denotes the interaction among its periphery nodes, and $n_{X}$ denotes the number of nodes in group $X$. We can now substitute the corresponding terms of Eq. (2) into Eq. (1), which gives us
@@
$P_{AEI}$ $=\overbrace{(\underbrace{\widehat{i}_{c_{A}}}_{\text{elite cohesion}}+\underbrace{\widehat{i}_{cp_{A}}}_{\text{mass amplification}}+\underbrace{\widehat{i}_{p_{A}}}_{\text{mass cohesion}})}^{\text{Group A's contribution}}$
$\quad+\underbrace{(\,\widehat{i}_{c_{B}}+\widehat{i}_{cp_{B}}+\widehat{i}_{p_{B}})}_{\text{Group B's contribution}}-\underbrace{2\,\widehat{e}_{AB}}_{\text{bridge}},$ (3)
where we use the hat symbol to denote the normalized form, where each component is divided by the denominator in Eq. 1.
@@
We refer to Eq. (3) as polarization decomposition as it allows us to further study the direct effects of different groups and hierarchies on the polarization score. By splitting the observed polarization score into components, we can detect the differences between groups and their hierarchies on the polarization measure and determine whether their contributions have changed over time. When we say ‘contribution’ or ‘impact’ of a certain group on observed polarization, we specifically refer to these components. This is valuable information in the context of polarization, where we are often interested in monitoring the evolution of the opposing groups and their power dynamics.
@@
For separating the total contribution of groups from the expected contribution of an individual member of the group, we also measure the marginal polarization for both core and periphery nodes. We define marginal polarization as the change in structural polarization when a node of a specific type is added to the network, which enables us to estimate how introducing an average elite user versus an average non-elite user affects the network’s overall polarization.
@@
A core node in group $A$ has $k_{c_{A}}$ links to other core nodes in the same group, it receives $k_{cp_{A}}$ links from the periphery and forms $k_{out}$ external links to group $B$. For a periphery node belonging to group $A$, it can have $k_{p_{A}}$ links to other periphery nodes. It amplifies $k_{cp_{A}}$ core nodes and connects to $k_{out}$ links outside its own group. By substituting the average of observed values into these decomposed degrees, we can approximate the marginal polarization for a mean core node as
@@
$\Delta_{c_{A}}P_{AEI}\approx\frac{2}{\alpha}\left(\frac{\langle k_{c_{A}}\rangle+\langle k_{cp_{A}}\rangle}{n_{A}^{2}}-\frac{\langle k_{out_{B}}\rangle}{n_{A}n_{B}}\right)\,,$ (4)
@@
where $\alpha$ denotes the denominator of Eq. 1. The formula is equivalent for the situation where an average periphery node is added to the polarized system $(\Delta_{p_{A}}P_{AEI}$), but you need to substitute the $k$ values with those reflecting the link distribution of peripheral nodes. Interestingly, the formula suggests that having an equal number of links in and out is not enough to have zero net impact on polarization when group $A$ is smaller than group $B$. See Appendix A for full derivation of this marginal polarization formula.
@@
In addition to structural polarization, we aim to quantify the degree of issue alignment separately for the elites and masses. Issue alignment contributes to political polarization by creating a situation where individuals or groups align themselves strictly along specific issues. When people strongly identify with a particular issue or group and use it as a defining factor for their political stance towards other topics, it can lead to several detrimental consequences [10]. We measure alignment using the normalized mutual information (NMI) method [10] for cores and peripheries separately. In essence, it quantifies the extent to which knowing the group of an user in one topic informs us about the users’s group in another topic.
@@
For estimating each user’s stance on a specfic issue, we use their group membership, $A_{1}$ or $B_{1}$ for the first network and $A_{2}$ or $B_{2}$ for the second network. Since this assessment requires user’s presence in both networks, we add the stances for each node in $V_{1}\cap V_{2}$ in stance vectors, $s_{1}$ and $s_{2}$, corresponding to the respective issues, i.e. these stance vectors encode each user’s group membership in the networks. Finally, we compute the issue alignment separately for the distinct hierarchies; NMI($s_{1}^{c}$, $s_{2}^{c}$) for the cores (elite alignment) and NMI($s_{1}^{p}$, $s_{2}^{p}$) for the peripheries (mass alignment). This measure is constrained between 0 and 1, where 0 indicates no alignment, and 1 indicates complete alignment between the issues in question.
@@ 4 Hierarchical Polarization in Finnish Parliamentary Elections 2019 and 2023
Figure 2: (A) demonstrates the increase in overall structural polarization across all networks, as measured by the AEI. Black bar corresponds to the proportion explained by the null model. The largest increase in the portion not explained by the null model was observed in the network representing economy-related discussions online. (B) The heatmaps depict the evolution of issue alignment over the four years, with 2019 on the left and 2023 on the right. Every pair of topics has experienced a substantial increase in the degree of alignment, as measured by the adjusted NMI. Climate and immigration were already reasonably aligned in 2019, however, the alignment doubled after four years. (C) illustrates the relationship between observed alignment and the average structural polarization scores for all topic pairs in both years—a slight linear relationship appears between the two quantities, possibly indicating that these phenomena are strenghtening each other. Note that in 2019, although some networks had high structural polarization scores, issue alignment remained relatively low. In contrast, by 2023, networks showed both high structural polarization and high issue alignment.
In this section, we first present the data for our case study, which investigates polarization trends during the Finnish Parliamentary Elections of 2019 and 2023. We demonstrate how the overall structural polarization and issue alignment have increased over four years, and then decompose the polarization scores separately for the elites and the masses.
@@ 4.1 Data
We use data collected from Twitter (Public API v1 and v2) during the Finnish parliamentary elections in 2019 and 2023. To ensure consistency when comparing these two snapshots, we consider a 12-week period leading up to the election day, resulting in the following periods: from January 21, 2019, to April 14, 2019 for the first elections, and from January 9, 2023, to April 2, 2023 for the second elections. For each year, five networks are constructed from sets of keywords related to larger topics, such as immigration and climate change (see Appendix in [10] for more details). Before constructing the networks, retweets that contain keywords belonging to two or more topics are removed from the dataset, as they would produce a systematic bias, i.e. inflate the extent of alignment between networks. This provides, in a sense, a lower bound for the actual alignment between the networks.
@@
Each network can be viewed as a graph where each user contributing to the topic is assigned to one node. In this graph, an edge between two nodes represents a retweet, which can be deemed as an agreement or a shared point of view on a selected issue between the corresponding users [22, 10, 56]. The largest connected component is extracted for each network [22, 10, 56], with the subsequent removal of self-loops and parallel edges [56, 10]. An average graph consists of approximately 10,000 nodes and 30,000 edges (see Table in Appendix C for more details).
@@
We tested the assumptions stated in the previous section independently: To verify the presence of political figures within cores, we assessed the probability of a politician being found in the core compared to the periphery. We found that politicians who ran for elections are at least 3-4 times more likely to be in the core than in the periphery (see Appendix D for more details). To view the connections between the outer periphery and the core nodes as mass amplification, we confirmed this assumption by checking that the majority of retweets flow towards the core nodes (see Appendix E for more details). Thus, we find our assumptions reasonable.
4.2 Drastic Increase in Polarization
Figure 3: (A) Polarization decomposition for structural contributions of different groups and their hierarchies to AEI-score. The figure illustrates the predominant influence of elite cohesion ($\widehat{i}_{{c}_{A}}$ $\widehat{i}_{{c}_{B}}$), mass amplification ($\widehat{i}_{{cp}_{A}}$ $\widehat{i}_{{cp}_{B}}$), and mass cohesion ($\widehat{i}_{{p}_{A}}$ $\widehat{i}_{{p}_{B}}$) on the overall score. The green part of spectrum represents the impact of the bridge between the opposing entities ($2\times\widehat{e}_{AB}$). The contributions of different hierarchical members not only vary within individual networks but also across the distinct networks. The part of the spectrum that corresponds to the internal structures is shifted to the left by an amount equal to the cross-interactions. This enables us to read the unadjusted AEI score for each network directly from the figure. (B) Groups vary in their sizes, and mostly consists of the masses. Smaller groups can have a great impact on the observed polarization.
Before breaking down polarization by groups and hierarchies, we first report the overall structural polarization and issue alignment across the networks in each snapshot (see Fig. 2). We confirm that every network passes the assortativity test described in Section 3.
Structural polarization has increased in all the networks studied here. Over these four years, climate and economy experienced the highest increase, when taking the null model [56] into account. Not only have the opposing groups moved further apart in some individual topics, but they have also become more aligned as the observed alignment values experienced an upward swing in each pair of topics studied here. In other words, the tendency of the same individuals to be divided on other topics as well has become more pronounced. In 2023, immigration aligned the most with the remaining topics, scoring the highest with economy (NMI = 0.6). The least (but substantially) aligned topics appear to be climate and economy, whose issue alignment was already at relatively high level in 2019.
@@ 4.3 Groups have Unequal Impact on the Overall Structural Polarization
Figure 4: (A) Elites are consistenly more aligned than masses across all topic pairs. Elites became more aligned in 2023, together with a smaller increase in the alignment of mass opinion on various issues. To capture the uncertainty around the observed values, we bootstrapped 500 pairs of networks for each topic pair. Each bootstrap sample represents a subgraph of the original network, where the sizes of cores and peripheries are subject to random fluctuations. We do this by sampling the nodes to groups according to their original group probabilities. (B) Elites tend to have higher marginal polarization as well compared to the mass. In both years, adding a new elite member to the economy network had the greatest impact. A weighted average of both groups’ marginal values is applied to obtain a single value representing the hierarchy’s mean effect on AEI.
Beyond assortativity, we find that each polarized group contains a hierarchy that cannot be explained by the Erdős–Rényi model. We confirmed assortativity if the two-group model yielded a lower description length than the single-group model. Similarly, hierarchy within each group was confirmed if the core-periphery-model resulted in a lower description length compared to the null model (see Section 3).
We show the contributions and sizes of different components in Fig. 3, where separate decompositions are created for each year separately for all topics. Here, groups labeled A (in red) refer to left-leaning groups, while groups labeled B (in blue) refer to right-leaning groups, respectively (see Appendix F for labeling details). This allows us to track the changes in the dynamics of left-right groups and their embedded hierarchies in the polarized milieu online.
@@
The polarization decompositions also reveal the influence and evolution of distinct hierarchical groups. For instance, in the immigration network, the most polarized topic in both years, the elite cohesion of the right-leaning group had a relatively large impact on polarization. The same elite were also significantly more organized than those in the left-leaning group, although the elite cohesion of the latter increased in 2023. The opposite phenomenon was observed for economic policy network, where the left-leaning group’s elite cohesion was higher, while right-leaning group elite evolved into more organized structure in 2023. Note how, in every network’s observed polarization, the role of at least one elite group’s cohesion had increased during the four years; Right-leaning elite became more cohesive in four networks and left-leaning elite in three networks, indicating a stronger role of elite in certain topics. The most drastic swap in terms of contributions of a hierarchical groups was witnessed in the mass amplification patterns in climate network, where right-leaning public outrun the previously dominating left-leaning public. Although in the absolute terms, the publics’ interaction patterns tend to explain most of the observed polarization in all networks, the relative contribution (when compared to the sizes of these hierarchical groups) of the elite members is, on average, greater, which we see in Fig. 3B and later in Fig. 4B.
@@
Communication between polarized groups has mostly dropped or remained at the same level across all topics. This type of communication, labeled as bridge, is the only interaction that could structurally decrease polarization. Its share of the decomposition has shrunk in all networks except the immigration network.
@@
Three main findings from the Fig. 3 stand out: first, the increase in structural polarization does not follow uniform patterns, second, a larger group size does not always lead to a higher contribution to the observed polarization, and third, the elites play relatively higher role than the masses. We see, for instance, how right-leaning group compensated its smaller size in climate and education with more organized core and greater mass amplification in 2023, leading to substantially higher impact on polarization when compared to the impact of left-leaning group. The greater relative contribution of the elite members is generally evident across all networks when reflecting it to the size of their groups shown in Fig. 3B. We also confirm this by computing the marginal polarization separately for the elite and mass individuals in Fig. 4B.
4.4 Elites are Significantly More Aligned than the Masses